# Fast, Modular, and Differentiable Framework for Machine Learning-Enhanced Molecular Simulations

### v1

@@ Abstract
We present an end-to-end differentiable molecular simulation framework (DIMOS) for molecular dynamics and Monte Carlo simulations. DIMOS easily integrates machine-learning-based interatomic potentials and implements classical force fields including particle-mesh Ewald electrostatics. Thanks to its modularity, both classical and machine-learning-based approaches can be easily combined into a hybrid description of the system (ML/MM). By supporting key molecular dynamics features such as efficient neighborlists and constraint algorithms for larger time steps, the framework bridges the gap between hand-optimized simulation engines and the flexibility of a PyTorch implementation. The superior performance and the high versatility is probed in different benchmarks and applications, with speed-up factors of up to 170×170$\times$170 ×. The advantage of differentiability is demonstrated by an end-to-end optimization of the proposal distribution in a Markov Chain Monte Carlo simulation based on Hamiltonian Monte Carlo. Using these optimized simulation parameters a 3×3$\times$3 × acceleration is observed in comparison to ad-hoc chosen simulation parameters. The code is available at https://github.com/nec-research/DIMOS.
@@
Molecular simulations are a cornerstone of modern computational physics, chemistry and biology, enabling researchers to understand complex properties of the system [1]. Traditional molecular dynamics (MD) and Markov Chain Monte Carlo (MCMC) simulations rely on pre-defined force fields and specialized software to achieve large timescales and efficient sampling of rugged free-energy landscapes [2]. However, conventional MD and MCMC simulation packages generally lack the flexibility and modularity to easily incorporate cutting-edge computational techniques such as machine learning (ML) based enhancements: Advances in machine learning interatomic potentials (MLIPs) promise improved accuracy for MD simulations [3], yet integrating these techniques into a scalable and user-friendly framework remains a major challenge, especially when developing novel approaches  [^1] For example, LAMMPS [38] and OpenMM [26] only recently introduced approaches to facilitate the use of MLIPs in simulations..
@@
Here we present an end-to-end differentiable molecular simulation framework (DIMOS) implemented in PyTorch [5], a popular library for ML research. DIMOS implements essential algorithms to perform MD and MCMC simulations, providing an easy-to-use way to interface MLIPs and an efficient implementation of classical force field components in addition to implementations of common integrators and barostats. Additional components are the efficient calculation of neighborlists and constraint algorithms which allow for larger timesteps of the numerical integrator. By relying on PyTorch, we inherit many advances achieved by the ML community: We achieve fast execution speed on diverse hardware platforms, combined with a simple-to-use and modular interface implemented in Python.
In particular, automatic differentiation, a cornerstone of ML, allows for novel approaches by enabling the automatic calculation of derivatives of arbitrary components in the computation graph. For example, differentiating through the simulation either to tune force-field parameters based on desired dynamic properties [6] or to adapt the parameters of simulation methods to accelerate the sampling in equilibrium [7] becomes possible, representing only a small subset of potential future developments.
@@
DIMOS fills an important void in the landscape of existing differentiable simulation frameworks which currently suffer from critical limitations. For example, torchMD [8] does not implement neighborlists and thus becomes computationally prohibitive for larger systems, and JAX M.D. [9] does not implement classical force field. While there exist separate implementations for many of the algorithms implemented in DIMOS in various code bases, there currently exists no unified and reliable simulation framework for them  [^2] For example, see https://github.com/openmm/NNPOps for implementations of neighborlists and particle-mesh Ewald methods..
In the remainder of the manuscript, we establish: a) The ease of use and modularity of the framework, b) the performance of the framework for molecular dynamics simulations, and c) showcase of the end-to-end optimization capabilities of the framework by optimizing Monte Carlo proposal distributions.
I Results
Figure 1: Computational Performance Analysis: Runtime $\tau$ in seconds per MD step for a water box with constant density as a function of number of atoms $N$. (a) Shows the runtimes for torchMD and DIMOS using single precision numerics and RFA to model the electrostatic interactions. (b) Demonstrates the influence of different approaches to model the electrostatic interactions: Runtimes are presented for RFA, Ewald, and PME for DIMOS using single precision numerical accuracy. In (c), we investigate the effect of the chosen computational precision on the runtime, modelling electrostatic interactions via RFA in DIMOS. In all cases, the lines indicate the theoretically expected asymptotic computational complexity. (d)-(f) Show the performance of running the MACE foundation model, where in (d) we compare the runtimes of DIMOS to ASE using the small MACE model and mixed precision for different $N$, and (e) investigates the influence of choosing different numerical precision within DIMOS. (f) Presents the runtimes for the biggest system that can be simulated using the large MACE model ($N=8844$) and compares the resulting $\tau$ for DIMOS and ASE.
@@ I Results
Setting up a minimal simulation in DIMOS can be done in only a few lines:
@@
Here, a molecular system configuration from a common input format is loaded and an integrator object to perform Langevin dynamics is initialized, which combined form the simulation; we then run the simulation for a given number of steps. All other required functions are hidden to the user, such as the handling of boundary conditions or neighborlists, but they are still easily accessible for inspection or extension. Due to the modular setup of the framework and since we are relying only on two external dependencies (PyTorch [5] to implement the simulation and parmed [11] to read configuration files), it is straightforward to develop novel approaches, a unique strength of our package.
In the following section, we present the resulting runtimes of DIMOS for simulations of water boxes to probe the influence of system size and simulation/modelling choices such as the method to calculate long-range interactions or different schemes for the numerical datatypes for classical force fields and MLIPs. Then we turn towards the simulation of proteins in explicit solvent, investigating the influence of constraints on the simulation speed and presenting ML/MM as hybrid solution. Finally, we demonstrate how the end-to-end differentiability of the code base allows us to differentiate through the whole simulation to tune parameters of the simulation and system in the framework of Hamiltonian Monte Carlo (HMC) [12, 13].
ml_atoms = [...] # Indices of atoms that are being modeled by ML system.
[…]
Figure 2: Simplicity of Use and Acceleration for Proteins: (a) Demonstration on how to set up both classically and machine-learning enhanced systems in DIMOS. ml_atoms| specifies which atoms should be modeled by the MLIP system, and also can be empty in case one does not want to model any systems using ML. (b-d) Achievable simulation time per day $\mathcal{T}$ with and without constrained motion of the hydrogens for each of the three cases of a system modeled purely by ML, the hybrid approach using ML/MM, and a purely MM based modelling. In all cases, the MLIP model was chosen to be mace-mp-large|. In (b), we present the data for alanine-dipeptide (Ala2), (c) shows Trp-Cage, and (d) shows dihydrofolate reductase (DHFR). All systems are modeled using explicit water modeled by the TIP3P parameters [14]. In (d), the biggest system, it is not possible to model purely with ML, as for this system size the MLIP model required more than the available GPU memory.
I.1 System-size Scalability
We investigate the runtimes per simulation step averaged over several simulation steps for water boxes of fixed density with varying number of atoms $N$. In Figs. 1(a)-(c), we present the average runtime per MD step $\tau$ as a function of number of atoms $N$ in an unconstrained simulation under Langevin dynamics at $300$ K with timestep $\Delta t=0.5$ fs, friction coefficient $\gamma=0.1$ fs-1, and cutoff $r_{c}=10$. The water is modeled by the TIP3P water model [14].
In (a) we compare $\tau$ obtained in DIMOS to torchMD [8] for simulations running in single floating point precision and electrostatic interactions modeled by the reaction field approximation (the RFA is the only setup for electrostatics supported by torchMD). We observe linear scaling of the runtime per step $t$ in DIMOS, whereas torchMD displays quadratic scaling as a function of number of atoms $N$. For the largest system size tested ($\approx 300,000$ atoms), DIMOS achieves a 170×170$\times$170 × speed-up over torchMD. This extrapolation is based on torchMD’s quadratic complexity, while we show that DIMOS maintains linear scaling due to neighborlists. For a fair comparison, we check the validity of the neighborlist after every integration step, avoiding any potential artifacts leading to inaccuracies in the force/energy evaluation.
Panel (b) demonstrates the influence of different methods how to model electrostatics within DIMOS. We use single precision numerics. Clearly the RFA is the fastest option due to the simple mean-field treatment of charges outside the cutoff. Somewhat slower, but still competitive is the particle-mesh Ewald (PME) [15] method, for which one expects $O(N\log N)$ complexity, which roughly captures our data. In contrast, using direct Ewald summation [16] with theoretically ideal complexity $O(N^{3/2})$ is significantly slower already for medium-sized systems, and can only provide a minor performance improvement over PME for small systems. In addition, direct Ewald summation requires significantly more GPU memory and can thus not be used for large systems.
The influence of the chosen numerical accuracy is presented in Fig. 1(c) with electrostatic interactions modeled by RFA, where we find that double precision as expected is overall slower on our employed GPU. Very little performance penalty is observed using mixed precision, where the integration is performed in double precision, but the force/energy evaluation is performed in single precision. However, this approach cannot be used for larger systems, as the memory requirements increase significantly. If memory consumption is not a bottleneck, it is advisable to always consider using mixed precision due to the reduction of accumulation of errors in comparison to single precision, where such effects can be detrimental.
In Figs. 1(d)-(f) we investigate the behavior of DIMOS in conjunction with MLIPs by simulating the same water boxes as in (a)-(c) with the mace-mp foundational model [17, 18]. MACE is a state-of-the-art equivariant message-passing graph neural network that models interatomic interactions. The simulation parameters remain largely the same, however, we now employ the standard interaction cutoff for message passing of $r_{c}=3$ in MACE. The runtimes are here compared to ASE [19], the de-facto (non-differentiable) standard approach in many situations for running MLIP based simulation due to its simplicity. We observe in (d) that for the smallest foundation model mace-mp-small, DIMOS displays a significant increase in speed using mixed precision (as also implicitly employed in ASE). As shown in (e), there is virtually no penalty in using mixed precision for simulations using MLIPs. This is because the runtime and memory consumption is dominated by the ML model, so that all other factors effectively become negligible, at least for the currently achievable system sizes.
Figure 1(f) shows the influence of choosing different sizes of the mace-mp model in comparison between DIMOS and ASE for the biggest system size ($N=8844$) that can be simulated using all ML model sizes. For the small model, the speed-up obtained by DIMOS is the biggest with factor 1.6×1.6$\times$1.6 ×, i.e., nearly half of the time in ASE is spent in integrating the equation of motion and memory transfer. For the bigger models, the difference between DIMOS and ASE become less pronounced due to the larger time spent in the actual ML model evaluation, but the speed-up still remains sizable.
Figure 3: End-to-end Optimization: Plots demonstrating the end-to-end learning of simulation parameters for HMC. (a) Loss $\mathcal{L}$ as function of training time $t$. (b),(c) The timestep $\Delta t$ and number of integration steps $L$ against training time $t$. (d),(e) The behavior of the masses $m$ for (d) the hydrogen and oxygen atoms in the water molecules, optimized as shared parameter between all molecules and (e) the remaining masses in Ala2. In (e) we plot $m-m_{0}$ to better highlight the relative change of masses for the diverse atom types. (f) Autocorrelation function $A(k)$ of the potential energy $U(x)$ as function of lag-time $k$ measured in terms of the number of energy evaluations.
I.2 Constraints and Hybrid ML/MM for Protein Systems
We simulate three benchmark systems of varying size: i) Alaninedipeptide (Ala2) with $2269$ atoms, ii) Trp-Cage with $7660$ atoms, and iii) dihydrofolate reductase (DHFR) with $23558$ atoms. The first two systems are standard test-systems to probe novel simulation methods, and the latter is a common testsystem from the joint AMBER/CHARMM benchmark  [^3] See https://ambermd.org/GPUPerformance.php.. Since we here also want to probe the effect of using constraint algorithms to increase the timestep on the achievable simulation time, we shift our metric to the simulation time achievable per wall-clock day $\mathcal{T}$, i.e., how many ns of simulation can be achieved per day, to evaluate the performance of DIMOS. We implement a combination of SETTLE [21] and CCMA [22] to constrain the relative position of all water molecules and all other bonds involving one hydrogen atom (for details, see Methods section III.5.4). In the simulations we use the Nosé–Hoover integrator with collision frequency of $0.05$ fs-1 and $\Delta t=0.5$ fs timestep for disabled constraints, and $\Delta t=2.0$ fs timestep with constraints enabled.
For the protein systems, we are considering three scenarios: The system purely modeled by mace-mp-large (ML), a hybrid approach where the protein is modeled by the MACE model (ML/MM), but the water is parameterized by the classical TIP3P water model [14], and a fully classical simulation using AMBER99SB [23] force field and TIP3P water (MM) [14]. In the ML and ML/MM case, electrostatic contributions are ignored (DIMOS implements several options), whereas in the MM case we use PME. As demonstrated in Fig. 2(a), it is straightforward to set up systems modeled by the three options in DIMOS. The only option to be specified by the user is which MLIP to use, and which atoms should be modeled by it.
In Fig. 2(b), we present the data for Ala2 using the three approaches with and without constraints. For the systems modeled by ML or ML/MM, we adopt the constraints as defined for the MM system. The ML approach is clearly the slowest option with $\mathcal{T}\approx 0.18$ ns/day of simulation time without constraints. This is in contrast to the fastest simulation time $\mathcal{T}\approx 20$ ns/day (more than 100×100$\times$100 × faster than the ML approach), achieved by modelling the interaction using MM and with constraints activated. As viable alternative, one may consider the hybrid approach, where with constraints we achieve $\mathcal{T}\approx 5.4$ ns/day.
For the larger Trp-Cage system presented in (c), the differences become even more pronounced. While a classical ML simulation without constraints achieves $\mathcal{T}\approx 0.05$ ns/day, the ML/MM approach without constraints is 14×14$\times$14 × faster. In the extreme case, employing constraints and modelling all interactions using MM, we observe a speed-up of >175×>175$\times$> 175 ×. While naturally modelling a part or all of the interactions using MM comes with an accuracy penalty, this highlights the strength of DIMOS as a modular framework with easy combination of various techniques tailored to the problem.
Figure 2(d) shows the performance for DHFR, which due to memory constraints of MACE could not be simulated completely using the ML approach. Using ML/MM, we obtain runtimes in the range of $\approx 0.15$ to $\mathcal{T}\approx 0.6$ ns/day for the ML/MM approach, and $\approx 3.8$ to $\approx 6.5$ ns/day for the pure MM modelling. This means that, in regions with increased accuracy requirements, it is now possible with DIMOS to simulate these systems with only a 10$\approx$ 10$\times$ 10 × performance penalty over purely MM modelling.
I.3 End-to-end Optimization of Hamiltonian Monte Carlo
We utilize the end-to-end differentiability of DIMOS to optimize the proposal distribution in MCMC simulations. HMC constructs proposal configurations that are informed by the forces acting on the particles, in contrast to random-walk MCMC where proposals are uninformed. This allows for large configurational changes without incurring a low acceptance rate. For more details on HMC and the learning setup, see Method section III.6.
The performance of HMC crucially depends on the choice of simulation parameters, where suboptimal choices can lead to a significant increase in autocorrelations between samples. We optimize three classes of parameters critical to the efficiency of HMC: The integration timestep $\Delta t$, the number of integration steps $L$, and the atomic masses $m_{i}$. The loss function $\mathcal{L}$ promotes large configurational changes while maintaining reasonable acceptance rates, balancing exploration and computational effort (for details, see Section III.6). Figures 3(a)-(e) present $\mathcal{L}$, $\Delta t$, $L$, and the atomic masses $m_{i}$ during training for a system of Ala2 in explicit solvent modeled by the classical force field AMBER99SB [23] and TIP3P water [14]. We clearly see that the loss $\mathcal{L}$ decreases as a function of training time $t$, and reaches a plateau towards the end. The timestep increases from the small starting value of to $\Delta t\approx 1$ fs, which is significantly larger than the $0.5$ fs timestep allowed by the physical dynamics in MD. In conjunction with the large number of integration steps of $L\approx 60$, this corresponds to big configurational changes for each proposal.
The masses of atoms in the water molecules are optimized for as shared parameters between all molecules. As shown in Fig. 3(d), we find that the hydrogen masses are increased while the masses of the heavy oxygen atoms are decreased, similar to what is done manually in MD simulations [24]. A similar trend can be observed in Fig. 3(e) for Ala2, however, with smaller relative changes in the atom masses.
In Fig. 3(f) we plot the autocorrelation function $A$ of the potential energy as function of lag time $k$ measured in the number of energy evaluations for a simulation where the parameters are fixed to the optimized ones. We find that, compared to a simulation with typical parameters of $\Delta t=0.5$ fs and $L=10$ without mass repartitioning shown in the same plot, the autocorrelations decay significantly faster in the optimized case. This translates to an exponential autocorrelation time (as measured in computational effort) of $\tau_{\mathrm{exp}}\approx 13,150$ in the trained case versus $\tau_{\mathrm{exp}}\approx 38,900$ in the case where the simulation parameters are chosen ad-hoc, corresponding to a 3$\approx$ 3$\times$ 3 × speed-up in sampling.
II Discussion
In this work, we have introduced DIMOS, an end-to-end differentiable simulation framework. Our results demonstrate that our framework not only extends the capabilities of conventional non-differentiable MD packages, but also provides a modular platform for implementing emerging methodologies in molecular simulations. Compared to prior approaches, our work bridges the gap between validated general-purpose simulation tools and the specialized needs of ML-enhanced modelling, allowing researchers to explore novel approaches that were previously computationally prohibitive.
We have shown that DIMOS has superior performance compared to other fully differentiable simulation frameworks, with a speed-up factor of up to 170×170$\times$170 ×. Ewald and particle-mesh Ewald summation are implemented as advanced approaches to calculate electrostatic interactions, allowing also for the gradient-based optimization of MLIPs based on them. The mixed precision implementation allows for accurate evolution in time without substantial overhead in computational effort.
Simulations using MLIPs are significantly accelerated using DIMOS, with speed-up of 1.6$\approx$ 1.6$\times$ 1.6 × over ASE. Another strength of DIMOS is demonstrated for three protein systems in explicit solvent, for which we implement a hybrid approach combining classical force fields and MLIPs. We find that this accelerates the simulations significantly over pure MLIP based modelling. Additional speed-up is achieved by constraining the movement of fast oscillating atoms, allowing for a 4×4$\times$4 × increased timestep compared to unconstrained dynamics. By optimizing end-to-end the parameters of HMC and introducing a data-driven way to perform mass-repartitioning, we are able to obtain a 3$\approx$ 3$\times$ 3 × acceleration in sampling efficiency over ad-hoc chosen simulation parameters.
DIMOS is designed to become a vital tool for researchers in computational physics, chemistry, and biology, combining the modularity needed for method development with good simulation performance. Due to the modular framework, it is straightforward to add additional features which may be needed in those communities, such as for example the support for more general simulation boxes or virtual particles.
III Methods
(a) Classical (MM)
(b) MLIP (ML)
(c) Hybrid approach (ML/MM)
(d) Typical energy components for a classical description of interactions based on the AMBER force field [23].
Figure 4: Schematic Visualization: In (a)-(c) a schematic representation of Alanine tripeptide (Ala3), where the atomic interactions are modeled (a) solely classically (MM), (b) solely using an MLIP (ML), or (c) using a hybrid approach (ML/MM). The red circle marks a region modeled by a classical force field, where one defines interaction terms based on covalent bonds and non-covalently bonded atoms. The blue circles represent the message passing iterations of a graph neural network used to construct many-body interactions, for details see Ref. [25] and the references therein. In all cases, only the operations based on calculating the force/energy for a single atom are depicted, and for a full evaluation the steps are repeated for every atom. In (d), the energy components for a classical force field are depicted, based on the parameterization of AMBER class interactions [23]. Covalently bonded and non-bonded interactions are distinguished, with exclusions in the non-bonded interactions for directly or indirectly bonded atoms. For more details on the individual contributions, see Section III.1.
We generally aim to reproduce the behavior of OpenMM [26] and check, whenever possible, consistency with regard to their implementation both for static and dynamic properties.
In the first part of the method section, we present the different approaches to model interatomic potentials in DIMOS. This first part will focus on classical force fields, in which a relatively simple and heuristic functional form is assumed to describe the interactions of the atoms, with explicit consideration of covalent bonds. In contrast, MLIPs typically do not consider explicit information of covalent bonds and instead often rely on learning the interactions based on generalized higher order terms. Finally, due to considerations of simulation speed, it may turn out to be beneficial for some systems to model part of the interactions classically, and another part using MLIPs. A schematic depiction of the three approaches is presented in Figs. 4(a)-(c), where for the depiction of the interaction via MLIPs a graph neural network with message-passing approach was considered.
The other parts of the Methods section present details related to the handling of the dynamical evolution of the system, such as integrators and constraints. Finally, we provide the details of the self-tuning HMC approach.
All simulations in the three benchmark simulation frameworks (DIMOS, OpenMM, ASE) for which we quote runtimes were performed using an NVIDIA RTX 6000 Ada GPU with 48GB memory on a server equipped with an Intel Xeon Gold 5412U CPU with 24 cores and 256GB memory.
III.1 Classical Force Fields

### v2

@@ Abstract
We present an end-to-end differentiable molecular simulation framework (DIMOS) for molecular dynamics and Monte Carlo simulations. DIMOS easily integrates machine-learning-based interatomic potentials and implements classical force fields including an efficient implementation of particle-mesh Ewald. Thanks to its modularity, both classical and machine-learning-based approaches can be easily combined into a hybrid description of the system (ML/MM). By supporting key molecular dynamics features such as efficient neighborlists and constraint algorithms for larger time steps, the framework makes steps in bridging the gap between hand-optimized simulation engines and the flexibility of a PyTorch implementation. We show that due to improved linear instead of quadratic scaling as function of system size DIMOS is able to obtain speed-up factors of up to $170\times$ for classical force field simulations against another fully differentiable simulation framework. The advantage of differentiability is demonstrated by an end-to-end optimization of the proposal distribution in a Markov Chain Monte Carlo simulation based on Hamiltonian Monte Carlo (HMC). Using these optimized simulation parameters a $3\times$ acceleration is observed in comparison to ad-hoc chosen simulation parameters. The code is available at https://github.com/nec-research/DIMOS.
I Introduction
@@
Molecular simulations are a cornerstone of modern computational physics, chemistry and biology, enabling researchers to understand complex properties of systems.[1] Traditional molecular dynamics (MD) and Markov Chain Monte Carlo (MCMC) simulations rely on pre-defined force fields and specialized software[2, 3] to achieve large timescales and efficient sampling of rugged free-energy landscapes.[4] These simulation packages generally lack the flexibility and modularity to easily incorporate cutting-edge computational techniques such as machine learning (ML) based enhancements without major changes to the code base: For example, the simulation engine LAMMPS[2] by now supports certain popular machine learning interatomic potentials (MLIPs)[5] such as MACE,[6] but novel architectures or new atomic descriptors require adaptation both of the model itself and the core simulation engine. Other ML based approaches relying on the end-to-end differentiability of the simulation framework are not at all supported by classical simulation packages. On the other hand, the design principle of ML focused packages is often prioritizing usablity over performance.
@@
Here, we present an end-to-end differentiable molecular simulation framework (DIMOS) implemented in PyTorch,[7] a popular library for ML research. Our design goal was a fast simulation engine in a high-level languange with ML based approaches at it’s core. For this, DIMOS implements essential algorithms to perform MD and MCMC simulations, providing an easy-to-use way to interface MLIPs and an efficient implementation of classical force field components in addition to common integrators and barostats. This is complimented by an implementation of efficient calculation of neighborlists and constraint algorithms which allow for larger timesteps of the numerical integrator. By relying on PyTorch, we inherit many advances achieved by the ML community: Fast execution speed on diverse hardware platforms, combined with a simple-to-use and modular interface implemented in Python.
@@
DIMOS is designed as a complementary engine for fast, differentiable prototyping and method development rather than a direct replacement for classical optimized MD codes, trading some peak performance for flexibility and end-to-end differentiability. The package fills an important void in the landscape of existing differentiable simulation frameworks[8, 9, 10, 11, 12] which currently suffer from respective limitations. For example the two PyTorch based packages we are aware of, torchMD[9] and TorchSim,[12] either do not implement neighborlists leading to computationally prohibitive simulation times or do not support setting up complex classical force field simulations. Note that TorchSim[12] was released after DIMOS. The other packages are implemented in either Jax or Julia: Jax M.D.[8] and FeNNol[11] are Jax[13] based, and Molly.jl[10] is implemented in Julia.[14] As consequence, the integration of PyTorch based code or models is not straightforward. Generally we want to note that while there exist separate implementations for many of the algorithms of DIMOS in various code bases, we are not aware of a unified simulation framework implementing them as comprehensively. [^1] For example, see https://github.com/openmm/NNPOps for implementations of neighborlists and particle-mesh Ewald methods.
Automatic differentiation, a cornerstone of ML, opens the door for novel approaches by enabling the automatic calculation of derivatives of arbitrary components in the computation graph. For example, differentiating through the simulation either to tune force-field parameters based on desired dynamic properties[16] or to adapt the parameters of simulation methods to accelerate the sampling in equilibrium[17] becomes possible, representing only a small subset of potential future developments.
In the remainder of the manuscript, we will discuss the components implemented in DIMOS, and how they can be interfaced in Section II. Section III presents benchmark results for water boxes highlighting the scaling behavior and the performance for protein systems is presented, discussing the use of hybrid classical and ML based modelling of interactions. The end-to-end differentiability is demonstrated in Section IV by tuning parameters of HMC for alanine-dipeptide in explicit solvent. Finally, we conclude in Section V.
(a) Classical (MM)
(b) MLIP (ML)
(c) Hybrid approach (ML/MM)
(d) Typical energy components for a classical description of interactions based on the AMBER force field.[18]
Figure 1: In (a)-(c) a schematic representation of Alanine tripeptide (Ala3), where the atomic interactions are modeled (a) solely classically ($M$), (b) solely using an MLIP (ML), or (c) using a hybrid approach (ML/$M$). The red circle marks a region modeled by a classical force field, where one defines interaction terms based on covalent bonds and non-covalently bonded atoms. The blue circles represent the message passing iterations of a graph neural network used to construct many-body interactions, for details see Ref. 19 and the references therein. In all cases, only the operations based on calculating the force/energy for a single atom are depicted, and for a full evaluation the steps are repeated for every atom. In (d), the energy components for a classical force field are depicted, based on the parameterization of AMBER class interactions.[18] Covalently bonded and non-bonded interactions are distinguished, with exclusions in the non-bonded interactions for directly or indirectly bonded atoms. For more details on the individual contributions, see Section II.1.
II Simulation Package
@@
Setting up a minimal simulation based on classical force fields in DIMOS can be done in only a few lines:
@@
Here, a molecular system configuration from a common input format (config.prmtop) is loaded and an integrator object to perform Langevin dynamics is initialized, which combined form the simulation; we then run the simulation for a given number of steps. All other required functions are hidden to the user, such as the handling of boundary conditions or neighborlists, but they are still easily accessible for inspection or extension. Due to the modular setup of the framework and since we are relying only on two external dependencies (PyTorch[7] to implement the simulation and parmed[20] to read configuration files), it is straightforward to develop novel approaches, which we believe is a strength of our package.
In the remainder of this section, we compactly present the simulation components implemented in DIMOS, and discuss their use and Python interface. For all components we implement, we aim to reproduce the simulation behavior of OpenMM[3] and have checked, whenever possible, consistency with regard to their implementation both for static and dynamic properties (corresponding test scripts are part of the published code base). Stochastic algorithms, such as the Langevin integrator, are implemented in a reproducible way by passing the random number generator object to every function relying on it.
User-facing examples and the scripts to generate the data used in the manuscript are provided with the codebase, and the docomentation includes a guide for integrating custom MLIP models into DIMOS’ PyTorch interface.
II.1 Classical Force Fields